A 20 kg brick stands upright on the ground as shown. What is the pressure it exerts on the ground?

Specific of brick : Height ( 10cm) Length (20cm) Width (5cm)


1
Expert's answer
2022-04-13T09:21:43-0400

The pressure exerted by a brick on the ground


"p=\frac{F}{A}=\frac{mg}{A}""p=\frac{20*9.8}{0.2*0.05}=19600\:\rm Pa"